OpenStudy (anonymous):

When using the substitution method, what expression for x in the first equation would be substituted in the second equation? x + 3y = –7 2x + 4y = –4 Choices: A.3y – 7 B.3y + 7 C.–3y – 7 D.–3y + 7

OpenStudy (anonymous):

2x+6y =-14, x= -7-6y =c

OpenStudy (anonymous):

x + 3y = –7 x = -3y-7 is C

OpenStudy (anonymous):

x + 3y = –7 2x + 4y = –4 x=-7-3y 2(-7-3y)+4y=-4 -14-6y+4y=-4 -2y=10 y=-5 x=-7-3(-5)=-7+15=8